# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part II: posterior mean approximation

### v2

@@ Appendix C Examples
By [4, Theorem 10.1], the solution $u$ of the heat equation above lies in $C((0,T];H^{1}_{0}((0,1)))$, and in fact $u(\cdot,t)$ has infinitely many continuous derivatives for each $t\in(0,T]$. By [25, Section 4.1], the solution can be written as $t\mapsto\exp(t\Delta)x^{\dagger}$. Let us define the linear map $g_{i}:\mathcal{H}\rightarrow\mathbb{R}$ by $g_{i}(h)=(\exp(t_{i}\Delta)h)(x_{i})$ for each $i$. Since $g_{i}$ is the composition of the linear and continuous maps $u\mapsto u(\cdot,t_{i})$, $C((0,T];H^{1}_{0}((0,1)))\rightarrow H^{1}_{0}((0,1))$ and $f\mapsto f(x_{i})$, $H^{1}_{0}((0,1))\rightarrow\mathbb{R}$, it follows that $g_{i}$ is linear and continuous. Then, with $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ defined by $Gh\coloneqq(g_{i}h)_{i=1}^{n}$, and with $\zeta\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}})$ where $\mathcal{C}_{\textup{obs}}=I$, this inverse problem is of the form (1).

### v3

@@ Appendix C Examples
By [5, Theorem 10.1], the solution $u$ of the heat equation above lies in $C((0,T];H^{1}_{0}((0,1)))$, and in fact $u(\cdot,t)$ has infinitely many continuous derivatives for each $t\in(0,T]$. By [30, Section 4.1], the solution can be written as $t\mapsto\exp(t\Delta)x^{\dagger}$. Let us define the linear map $g_{i}:\mathcal{H}\rightarrow\mathbb{R}$ by $g_{i}(h)=(\exp(t_{i}\Delta)h)(x_{i})$ for each $i$. Since $g_{i}$ is the composition of the linear and continuous maps $u\mapsto u(\cdot,t_{i})$, $C((0,T];H^{1}_{0}((0,1)))\rightarrow H^{1}_{0}((0,1))$ and $f\mapsto f(x_{i})$, $H^{1}_{0}((0,1))\rightarrow\mathbb{R}$, it follows that $g_{i}$ is linear and continuous. Then, with $G\in\mathcal{B}(\mathcal{H},\mathbb{R}^{n})$ defined by $Gh\coloneqq(g_{i}h)_{i=1}^{n}$, and with $\zeta\sim\mathcal{N}(0,\mathcal{C}_{\textup{obs}})$ where $\mathcal{C}_{\textup{obs}}=I$, this inverse problem is of the form (1).